OpenStudy (anonymous):

factor this quadratic, x^2-10x+9

OpenStudy (anonymous):

-x-9x=-10x -9(-1)=9 so we can factor it like this: (x-9)(x-1)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

but now how do u do this? x^2+3x-10

OpenStudy (anonymous):

5x-2x=3x 5(-2)=-10 so we can factor like this: (x+5)(x-2) make sense?

OpenStudy (anonymous):

hmm okay i know ur right, but im just not understanding your first step? how do u get 5x-2x=3x sorry im new to this class lol

OpenStudy (anonymous):

Well what you have to do is figure out what will work for you to end up with both the 3x and the -10. Since -2x+5x (or 5x-2x) equals 3x, and if you multiply -2 by 5 it gives you -10, those are the values to use when factoring it. Sorry, I'm not very good at explaining :)

OpenStudy (anonymous):

ehh sort of

OpenStudy (anonymous):

x^2+4x-12 what about this one? O.o

OpenStudy (anonymous):

Think about your factor of twelve: 2, 3, 4, 6... It's kind of like process of elimination. 4*3 is 12, but 4-3 doesn't equal four, nor does 3-4. So try something else. 2*6=12, 2-6 doesn't equal four, but 6-2 does! So we need two x's since it's x^2 (x )(x ) Since 2*6 is 12, these number go on the outside. (x 2)(x 6) Since 12 is negative in this problem, either 2 or 6 needs to be negative (since a negative*positive=negative) (x-2)(x+6) If we work it out: x*x+6x-2x-12 x^2+4x-12 we see that it works!

OpenStudy (anonymous):

factors*

OpenStudy (anonymous):

Hopefully that's a better explanation...

OpenStudy (anonymous):

yeah thanks :)

OpenStudy (anonymous):

x^2-2-12 this one i dont get i put (x-3)(x-4) its incorrect apparently

OpenStudy (anonymous):

Is that supposed to say 2x? Like x^2-2x-12?

OpenStudy (anonymous):

no its saying to the 2nd power

OpenStudy (anonymous):

Oh, so is it \[x^2-12\]

OpenStudy (anonymous):

its x^2-2x-12

OpenStudy (anonymous):

Oh. Hmmm... That's impossible to factor. Is that one of your options?

OpenStudy (anonymous):

thats the eqiation

OpenStudy (anonymous):

oh crap im sorry, its x^2-x-12

OpenStudy (anonymous):

Oh, that makes more sense :) Your answer is so close! But remember to get a negative number, like -12, one of the outside number needs to be positive and the other negative. So make it (x-3)(x+4) and that should be your answer :)

OpenStudy (anonymous):

Actually wait. In this one, order matter. 4 has to be the negative one. (x+3)(x-4) is your answer.

OpenStudy (anonymous):

matters* See, this is the importance of checking your answers afterward lol

OpenStudy (anonymous):

x*x-4x+3x-12 x^2-x-12

OpenStudy (anonymous):

thanks thats right, :) so then how would we do this one 10x^2+31x+15

OpenStudy (anonymous):

Okay, so let's start with 10. The only factors of 10 are 5 and 2 (and technically, 1, but we're going to forget about that). So, 2x and 5x will be put on the left since 2x*5x=10x^2 (5x )(2x ) Now, let's look at 15. The only factors of 15 are 3 and 5 (forgetting about 1 again). So, 3 and 5 need to go on the outside so we can achieve the "+15" part of the equation. (5x 3)(2x 5) Now, for 31. 31 is positive, so (5x+3)(2x+5) As always, we need to check it to make sure everything is in the right order :) As luck would have it, this works out without us having to really figure out how to get 31. 5x*2x+25x+6x+15 10x^2+31x+15

OpenStudy (anonymous):

Sorry, but I have to go. If you have other struggles, bump your question or create a new one and maybe someone else can help you :) Good luck!

OpenStudy (anonymous):

okay thank u girl!